Previous

Calendar

March

2006

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

26

JD57

27

JD58

28

JD59

1

JD60
Site Visit (KR, MW): Details

Soils Sampled (MW)

Laser LN2 filled from 18 to 27cm. PFP swap.

2

JD61 (warm/calm)

3

JD62

4

JD63

5

JD64
Site Visit (KR): Details

Laser LN2 filled from 15 to 27.5cm.

6

JD65 (warm/calm)
Site Visit (SpB): Details , Photos (35)

from 11:25-11:30 MST, swiped krypton hygrometer

from 12:15-13:01 MST, Removed LI-6262, now using IRG3-224

from 13:20-13:35 MST, Changed profiler (li-6251) dessicant

4-flask swap at East/West Robots

7

JD66

8

JD67
Site Visit (KR): Details

Laser LN2 filled from 19 to 30cm.

9

JD68

10

JD69

11

JD70
Site Visit (KR): Details

Laser LN2 filled from 22 to 34cm.

12

JD71

13

JD72

14

JD73
(Full "Worm" Moon)

15

JD74
Site Visit (MW): Details

Soils Sampled (MW)

16

JD75

17

JD76

18

JD77

19

JD78
Site Visit (KR): Details

Laser LN2 filled from 20 to 40cm.

20

JD79 (snow)
Spring Equinox

21

JD80 (snow)

22

JD81
Site Visit (KR): Details

Laser LN2 filled from 27cm to 40cm. N2 (ref) flow rate set to 8psi.

23

JD82
spectrometer study (NM)

24

JD83 (warm)
Site Visit (SpB, DM, AW, NM, D): Details , Photos (26)

Helicopter fly-by

at 10:30 MST, swiped krypton hygrometer

from 12:50-13:30 MST, Added check-valve to the li-6262 ec

at ~14:30 MST, swapped span gas (old cylinder=CC156367, new cylinder= CC86948, at 1800 psi)

4-flask swap at East/West Robots

Laser LN2 filled from ~35cm to FULL (approx 40cm?).

spectrometer study (NM)

25

JD84
spectrometer study (NM)

26

JD85

27

JD86

28

JD87

29

JD88

30

JD89

31

JD90
CU Holiday (Spring Break)

Site Visit (JH, MP, AT): Details

slow-response ozone instrument turned on (AT)

1

JD91